# Nellie: A Neuro-Symbolic Inference Engine for Grounded, Compositional, and Explainable Reasoning

### v4

@@ 4.1 Inference Rule Structure
Via template I, the system proves the hypothesis by finding a provenance Fact stored in its knowledge store that entails the hypothesis. Via template II, it enumerates a pair, Fact1 and Fact2, both either stored in the knowledge store or themselves recursively proved, such that the pair in conjunction entail the hypothesis.  [^4] Note that a strict two-premise decomposition can prove expressively powerful. For example, to prove a hypothesis such as those in EntailmentBank (Dalvi et al., 2021) that require a three-premise conjunction $h\Leftarrow f_{1}\land f_{2}\land f_{3}$, Nellie produces instead a recursive set of decompositions $H\Leftarrow f_{1}\land f_{i};f_{i}\Leftarrow f_{2}\land f_{3}$. Template I is given higher search precedence than II, yielding an intuitive high-level procedure: we first look up the hypothesis against the factbase, searching for an entailing fact. If we do not find one, we decomposes the hypothesis into a pair of statements to be proved. Concretely, for an input hypothesis $h$, we define the predicate $\textsc{Prove}(h)$ that serves as the primary goal term. We define the following core meta-rules, which use the neural predicates Retrieve, Entails, and RuleGen. At each step in the backward-chaining search, Nellie’s Prolog engine attempts unify a query with the head of one of these three rules:
@@
    $\textsc{Prove}(h)\Leftarrow\textsc{Retrieve}(h^{+},f_{1}^{-})\ \land\ \textsc{RuleGen}(h^{+},f_{1}^{+},f_{2}^{-})$ $\quad\land\textsc{Entails}([f_{1},f_{2}],h)\ \land$ $\textsc{Prove}(f_{2})$
@@ 4.2 Unification with Retrieved Facts
For factbase fact $f$, $\textsc{Prove}(f)$ is vacuously true. Rule 1 shows how we prove $\textsc{Prove}(h)$ using retrieval. The predicate Retrieve proposes candidate $f_{i}$’s given $h$ using a FAISS (Johnson et al., 2019)-based nearest neighbor dense retrieval index. [^5] For some situational questions, it is necessary to ground a subquery in evidence from the problem, e.g. “A ball is rolling down a hill.” To handle this, we add the question setup (defined as all but its last sentence) as a “fact” always proposed by Retrieve. We train the retrieval encoder via ranking loss such that the embedding for a hypothesis is maximally similar to its supporting facts. To promote logical coherence and improve the precision of the system, we apply a set of differently-trained neural RTE models as filters $\textsc{Entails}_{j}(\cdot)$ that iteratively rule out $f_{i}$ candidates that are not classified as entailing $h$.
@@ 4.3 Dynamic Rule Generator (DRG)
If we do not find an entailing fact for hypothesis $h$, we to decompose $h$ into a pair of entailing premises; we propose candidates using nucleus sampling (Holtzman et al., 2019) from a seq2seq model. The predicate $\textsc{RuleGen}(h,f_{1},f_{2})$ prompts a model trained to generate $h\rightarrow f_{1},f_{2}$ pairs. The space of potential decompositions is very large; to bias the proof search towards those more likely to be grounded in the corpus, we adopt a two-pronged approach, illustrated in inference rules 2 and 3, that proposes a heterogeneous search frontier using different biasing strategies.
@@
Template Conditioned Generation One way in which we might improve our LM-based proposal function is to leverage the high-level structure that supports reasoning in a given domain through template-guided generation, in this case to bias search towards the semi-structure of WorldTree. [^6] While our approach is applicable to a wide array of reasoning problems, the use of WorldTree-specific templates illustrates one way by which we can infuse domain-specific structure into the neural search algorithm. We note that this is a strict departure from the burdensome process of symbolic knowledge engineering; it is instead a simple means by which we can enlist a subject matter expert in order to improve the effectiveness of the engine. As our experiments show, this module improves Nellie’s QA accuracy by a few points, though ablating it yields a similarly strong performance. WorldTree tables correspond to types of facts with similar syntax and semantics that support scientific reasoning. WorldTree QA items are annotated with the facts that support their answers. Each table can be viewed as an n-ary relation of columns whose values are text spans. E.g., the Taxonomic relation has columns <A>, [HYPONYM], <is a kind of>, <SCOPE1>, [HYPERNYM], <for>, [PURPOSE]. Rows include ‘a bird is an animal’ and ‘a seed is a kind of food for birds.’
@@
Accordingly, for the RuleGen predicate in rule 2, half of the $f_{1},f_{2}$ candidates are sampled from the DRG conditioned on $h$ plus a template that cues the model to reflect the syntax of a WorldTree table. We train the DRG to accept any masked infilling template (e.g. <mask>is a kind of <mask>, akin to those used to pretrain recent LMs (Lewis et al., 2019; Raffel et al., 2020)), and propose decompositions whose first fact reflects the template’s syntax. We thus create a $h,t_{1}\rightarrow f_{1},f_{2}$ model. We feed the model templates drawn from WorldTree’s tables, guiding it towards proof steps more likely to be grounded in the factbase. We reuse the same model for non-template-conditioned generation by feeding it an empty template. In practice, we make two generation calls: one samples $m$ free-generated candidates, and a second samples $n$ candidates for each of $n_{t}$ templates (see §C).
@@
Template Selection WorldTree is a diverse corpus, containing tables that are specific to a particular subset of science problems (e.g. the “Predator-Prey” table). As conditioning on dozens of templates can be computationally expensive, we introduce a case-based reasoning (Schank, 1983; Das et al., 2021) approach that selects relevant templates for a given hypothesis. We construct an Okapi-BM25 (Jones et al., 2000) retrieval index over questions from the WorldTree QA training set to obtain the most lexically similar items to a query. At inference time, we select the top-$k$ most similar questions to the query, and take as our template subset the tables of the questions’ annotated facts.
@@
Retrieval Conditioned Generation In rule 3, rather than generate a pair of subqueries, we immediately ground half of the antecedent by choosing as $f_{1}$ a fact retrieved directly from the corpus. We have the DRG force decode $f_{1}$ before generating $f_{2}$ as normal, then recur only on $f_{2}$.
@@
Filters As stochastic sampling from LMs can be a noisy process, some fraction of generated candidates may be invalid: generated facts may be incoherent or not compositionally entail $h$. Accordingly, we introduce a suite of compositional entailment verifiers (Khot et al., 2020; Jhamtani and Clark, 2020) trained on two-premise compositional entailment. We also add a “self-ask” filter, which following (Tafjord et al., 2022) is an LM fine-tuned to assign a statement a truth value. If the confidence is below 0.5 for either an entailment judgement or the ‘self-ask’ belief in $f_{1}$ or $f_{2}$, then we filter the pair. All filters condition on the question setup text. When Nellie uses these rules, the unification score equals the lowest of the scores for $\textsc{Prove}(f_{1})$, for $\textsc{Prove}(f_{2})$, and the confidence of an entailment filter $s_{e}([f_{1},f_{2}]\Rightarrow H)$.
@@ 4.4 Proof Search
Given a query, Nellie searches for $t$ seconds to find up to $p$ proofs of depth $d$ or less. We follow (Weber et al., 2019) in pruning search branches whose unification score is guaranteed to fall below the current best, given our monotonic aggregation function $\min(\cdot)$. Found proofs that score under the current best do not count towards $p$. Full pseudocode for the algorithm, which follows a depth-first search with a breadth-first lookahead (Stern et al., 2010) to check for unification of generated subgoals, can be found in §E.
@@ 5 Experiments
Our experiments illustrate how the approach exemplified by Nellie provides grounded and logical explanations, performing comparably or better than approaches that do not satisfy these properties. We evaluate models on two multiple-choice QA datasets constructed so that correct answers are supported by facts in the WorldTree corpus: EntailmentBank (Dalvi et al., 2021) is a dataset of entailment trees for declarativized ARC answers, showing how the hypothesis can be derived via compositional entailment hops from WorldTree facts. We recast the test set, initially designed to test tree reconstruction, into QA by retrieving the corresponding multiple-choice ARC questions from which the hypotheses were constructed. WorldTree (Xie et al., 2020) is a subset of the ARC dataset (Clark et al., 2018) annotated with undirected explanation graphs whose nodes are facts from the WorldTree tablestore. We note that WorldTree explanations do not show how facts should combine. There is no guarantee that fully grounded trees exist for these questions using the WorldTree corpus alone. The difficulty of this task is akin to that of a course exam: the teacher (us) provides the student (the model) a long study guide of facts, but expects the student to use these facts via novel composition in order to reason coherently about a problem.
Our task metric is accuracy: whether a generated proof of the correct option outscores any other. [^7] If it produces no proofs, it gives no answer and is wrong. Nellie searches for up to $p$=10 proofs of max depth $d$=5 with a timeout of $t$=180 seconds per option. §B discusses further hyperparameters.
@@
Baselines We evaluate Nellie against Entailer (Tafjord et al., 2022), another system that produces entailment tree proofs via backchaining. Entailer stops recurring not when it finds entailing facts from a corpus, but rather when the model believes that a subquery is true with high confidence. Its trees are thus not grounded in verified facts. We reimplemented their algorithm using their T5-11B-based model, reporting their configuration of a max tree depth (D) of 3 and minimum of 1 (i.e. $\geq 1$ decomposition). We also evaluate their ablated setting which recurs exactly once (D=1). This is a baseline that is neither grounded nor particularly interpretable, generating just a pair of statements. To isolate the impact of model size, we also evaluate an Entailer-3B model based on the same T5-3B model as Nellie’s DRG. [^8] We obtained the training data from (Tafjord et al., 2022) and trained our model in consultation with the authors.
We also compare against grounded xQA methods without logical structure (see Fig 2): PathNet (Kundu et al., 2019), which constructs 2-hop chains by linking entities between facts, and three approaches that build graphs from facts using linear programming: TupleILP (Khot et al., 2017), ExplanationLP (Thayaparan et al., 2021), and Diff-Explainer (Thayaparan et al., 2022). [^9] We show numbers reported by (Thayaparan et al., 2021, 2022).
Figure 4: Effect of ablating one or both of rule-conditioned (RCG) and template-conditioned generation (TCG)
@@ 5.1 Results
Figure 4 shows the results of ablating Nellie’s two conditional generation modules and replacing them with vanilla generation. Removing both template- and retrieval-conditioned generation lowers Nellie’s performance in all circumstances, highlighting the empirical benefit of structured guidance. Ablating TCG (Nellie row vs. (-- TCG) and (-- RCG) vs. (-- Both)) drops performance by 1-1.5 points. Ablating RCG (Nellie vs (-- RCG)) drops it by 3-4.
@@ 5.2 Domain Generalization and Knowledge Scaling
While Nellie was trained on datasets centered around WorldTree, we show that it can perform in another domain by altering the knowledge over which it reasons. We consider OpenBookQA (Mihaylov et al., 2018), a dataset that requires reasoning over facts not included in WorldTree; each question is associated with one WorldTree science fact (“metals conduct electricity”), but also one fact from a separate pool of common knowledge (“a suit of armor is made of metal”).
As with a classical expert system, we have designed Nellie to be “improvable merely by making statements to it” (McCarthy, 1959). This means that as we increase its provided knowledge store, we should expect Nellie to reason more effectively. Because WorldTree does not cover all the knowledge required to answer OBQA questions, we test whether Nellie performs better on them when we add the common knowledge to its retrieval index. As the common knowledge annotations are one fact per question and are not designed specifically for entailment, we note that it is unlikely a priori that a fully grounded entailment tree can be created using the available knowledge alone, making this a challenging task for our reasoner.

### v5

@@ 4.1 Inference Rule Structure
Via template I, the system proves the hypothesis by finding a provenance Fact stored in its knowledge store that entails the hypothesis. Via template II, it enumerates a pair, Fact1 and Fact2, both either stored in the knowledge store or themselves recursively proved, such that the pair in conjunction entails the hypothesis. [^4] We find that two-premise decomposition is sufficiently powerful and expressive for our purposes. For example, to prove a hypothesis such as those in EntailmentBank Dalvi et al. (2021) that requires a three-premise conjunction $h\Leftarrow f_{1}\land f_{2}\land f_{3}$, Nellie produces instead a recursive set of decompositions $H\Leftarrow f_{1}\land f_{i};f_{i}\Leftarrow f_{2}\land f_{3}$. Template I is given higher search precedence than II, yielding an intuitive high-level procedure: we first look up the hypothesis against the factbase, searching for an entailing fact. If we do not find one, we decomposes the hypothesis into a pair of statements to be proved. Concretely, for an input hypothesis $h$, we define the predicate $\textsc{Prove}(h)$ that serves as the primary goal term. We define the following core meta-rules, which use the neural predicates Retrieve, Entails, and RuleGen. At each step in the backward-chaining search, Nellie’s Prolog engine attempts to unify a query with the head of one of these three rules:
@@
    $\textsc{Prove}(h)\Leftarrow\textsc{Retrieve}(h^{+},f_{1}^{-})\ \land\ \textsc{RuleGen}(h^{+},f_{1}^{+},f_{2}^{-})$
    $\quad\land\textsc{Entails}([f_{1},f_{2}],h)\ \land$ $\textsc{Prove}(f_{2})$
@@ 4.2 Unification with Retrieved Facts
For factbase fact $f$, $\textsc{Prove}(f)$ is vacuously true. Rule 1 shows how we prove $\textsc{Prove}(h)$ using retrieval. The predicate Retrieve proposes candidate $f_{i}$’s given $h$ using a FAISS Johnson et al. (2019)-based nearest neighbor dense retrieval index. We train the retrieval encoder via ranking loss such that the embedding for a hypothesis is maximally similar to its supporting facts. To promote logical coherence and improve the precision of the system, we apply a set of neural models for recognizing textual entailment (RTE) as filters $\textsc{Entails}_{j}(\cdot)$ that iteratively rule out $f_{i}$ candidates that are not classified as entailing $h$.
@@
For some questions such as the one depicted in Figure 4, it is necessary to ground a subquery in evidence from the problem. To handle this, we add the question setup (defined as all but its last sentence) as a “fact” always proposed by Retrieve.
@@ 4.3 Dynamic Rule Generator (DRG)
If we do not find an entailing fact for hypothesis $h$, we decompose $h$ into a pair of entailing premises; we propose candidates using nucleus sampling Holtzman et al. (2019) from a seq2seq model. The predicate $\textsc{RuleGen}(h,f_{1},f_{2})$ prompts a model trained to generate $h\rightarrow f_{1},f_{2}$ pairs. The space of potential decompositions is very large; there are many deductively valid ways to prove a hypothesis in natural language, though only a fraction of them are ultimately groundable in the provided factbase. To bias the proof search towards those more likely to be grounded in the corpus, we adopt a two-pronged approach, illustrated in inference rules 2 and 3, that proposes a heterogeneous search frontier using different biasing strategies in addition to straightforward LM sampling.
Figure 4: Example question in which a sub-hypothesis in the proof tree is grounded to the question context rather than to the factbase.
Template Conditioned Generation (TCG)
@@
One way in which we improve our LM-based proposal function is to leverage the high-level structure that supports reasoning in a given domain: we propose template-guided generation to bias search towards the semi-structure of WorldTree. [^5] While our approach is applicable to a wide array of reasoning problems, the use of WorldTree-specific templates illustrates one way by which we can infuse domain-specific structure into the neural search algorithm. This is a strict departure from the burdensome process of symbolic knowledge engineering. As our experiments show, this guidance method improves Nellie’s QA accuracy by a few points, though ablating it yields a similarly strong performance. WorldTree tables correspond to types of facts with similar syntax and semantics that support scientific reasoning. WorldTree questions are annotated with the fact rows that support their answers. Each table can be viewed as an n-ary relation of columns whose values are text spans. E.g., the Taxonomic relation has columns <A>, [HYPONYM], <is a / a kind of>, <SCOPE1>, [HYPERNYM], <for>, [PURPOSE]. Rows include ‘a bird is an animal’ and ‘a seed is a kind of food for birds.’
@@
Thus, for the RuleGen predicate in rule 2, half of the $f_{1},f_{2}$ candidates are sampled from the DRG conditioned on $h$ plus a template that cues the model to reflect the syntax of a WorldTree table. We train the DRG to accept any masked infilling template (e.g. <mask>is a kind of <mask>, akin to those used to pretrain LMs Lewis et al. (2019); Raffel et al. (2020)), and propose decompositions whose first fact reflects the template’s syntax. We thus create a $h,t_{1}\rightarrow f_{1},f_{2}$ model. We feed the model templates drawn from WorldTree’s tables, guiding it towards proof steps more likely to be grounded in the factbase. A sample of the 150 templates can be found in Figure 5 (a larger list is shown in §D). We reuse the same model for non-template-conditioned generation by feeding it an empty template. In practice, we make two generation calls: one samples $m$ free-generated candidates, and a second samples $n$ candidates for each of $n_{t}$ templates.
WorldTree Relation Template
KINDOF <mask>is a kind of <mask>
IFTHEN if <mask>then <mask>
PROP-THINGS <mask>has <mask>
CAUSE <mask>causes <mask>
MADEOF <mask>made of <mask>
REQUIRES <mask>requires <mask>
ACTION <mask>is when <mask>
USEDFOR <mask>used to <mask>
PARTOF <mask>a part of <mask>
CHANGE <mask>changes <mask>
USEDFOR <mask>used for <mask>
AFFECT <mask>has <mask>impact on <mask>
PROP-ANIMAL-ATTRIB <mask>is a <mask>animal
SOURCEOF <mask>is a source of <mask>
COMPARISON <mask>than <mask>
EXAMPLES an example of <mask>is <mask>
COUPLEDRELATIONSHIP as <mask>decreases <mask>will <mask>
PROP-CONDUCTIVITY <mask>is <mask>conductor
PROP-GENERIC <mask>is a property of <mask>
HABITAT <mask>live in <mask>
MEASUREMENTS <mask>is a measure of <mask>
Figure 5: Sample of WorldTree templates used for guided generation.
Template Selection
@@
WorldTree is a diverse corpus, containing tables that are specific to a particular subset of science problems (e.g. the “Predator-Prey” table). As conditioning on dozens of templates can be computationally expensive, we introduce a case-based reasoning Schank (1983); Das et al. (2021) approach that selects relevant templates for a given hypothesis. We construct an Okapi-BM25 Jones et al. (2000) retrieval index over questions from the WorldTree QA training set to obtain the most lexically similar items to a query. At inference time, we select the top-$k$ most similar questions to the query and take as our template subset the tables of the questions’ annotated facts.
Retrieval Conditioned Generation (RCG)
@@
In rule 3, rather than generate a pair of subqueries, we immediately ground half of the antecedent by choosing as $f_{1}$ a fact retrieved directly from the corpus. We have the DRG force decode $f_{1}$ before generating $f_{2}$ as normal, then recur only on $f_{2}$.
Filters
@@
As stochastic sampling from LMs can be noisy, some fraction of the generated candidate set may be invalid: premises may be incoherent, or the decomposition might not properly entail $h$. Accordingly, we introduce a set of compositional entailment verifiers Khot et al. (2020); Jhamtani and Clark (2020) trained on two-premise compositional entailment. We also add a “self-ask” filter, which following Tafjord et al. (2022) is an LM fine-tuned to assign a statement a truth value. If the confidence is below 0.5 for either an entailment judgment or the ‘self-ask’ belief in $f_{1}$ or $f_{2}$, then we filter the pair. All filters condition on the question text as context. When Nellie uses these rules, the unification score equals the lowest of the scores for $\textsc{Prove}(f_{1})$, for $\textsc{Prove}(f_{2})$, and the confidence of entailment filters $s_{e}([f_{1},f_{2}]\Rightarrow H)$.
@@ 4.4 Proof Search
Given a query, Nellie searches for $t$ seconds to find up to $p$ proofs of depth $d$ or less. We follow Weber et al. (2019) in pruning search branches whose unification score is guaranteed to fall below the current best, given our monotonic aggregation function $\min(\cdot)$. Found proofs that score under the current best do not count towards $p$. Full pseudocode for the algorithm, which follows a depth-first search with a breadth-first lookahead Stern et al. (2010) to check for the unification of generated subgoals, can be found in §E. It is parameterized by
  1. 1.
    A maximum number of proofs $m$ at which to cut off searching. In experiments, we set this to $10$ for top-level queries and $2$ for recursive subqueries.
  2. 2.
    A number of support facts $n_{f}$ to retrieve at each call to $\textsc{Retrieve}_{K}$, which we set to 15.
  3. 3.
    Candidate generation rates $n_{v}$ for vanilla nucleus-sampled decompositions, $n_{t}$ for template-conditioned decompositions, and $n_{r}$ for retrieval-conditioned generations. We set these each to $40$. [^6] Due to batching, these correspond to 3 total calls to the HuggingFace library’s generate function for the T5-3B model. Upon removing exact match duplicates, a call to RuleGen produces about $100$ candidates.
  4. 4.
    Entailment scoring module $\textsc{Score}_{e}(\cdot)$, which is a separate RTE cross-encoder model for single- and double-premise entailments.
@@ 5 Experiments
Our experiments illustrate how the approach exemplified by Nellie provides grounded and logical explanations, performing comparably or better than approaches that do not satisfy these properties. We evaluate models on two multiple-choice QA datasets constructed so that correct answers are supported by facts in the WorldTree corpus:
EntailmentBank Dalvi et al. (2021) is a dataset of entailment trees for declarativized answers to the AI2 Reasoning Challenge (ARC) dataset Clark et al. (2018), showing how the hypothesis can be derived via compositional entailment hops from WorldTree facts. We recast the test set, initially designed to test tree reconstruction, into QA by retrieving the corresponding multiple-choice ARC questions from which the hypotheses were constructed.
WorldTree Xie et al. (2020) is a subset of the ARC dataset annotated with undirected explanation graphs whose nodes are facts from the WorldTree tablestore. We note that WorldTree explanations do not show how facts should combine. There is no guarantee that fully grounded trees exist for these questions using the WorldTree corpus alone. The difficulty of this task is akin to that of a course exam: the teacher (us) provides the student (the model) with a very large study guide of facts, but expects the student to use these facts by composing them to reason coherently about a problem.
Our task metric is accuracy: whether a generated proof of the correct option outscores any other. [^7] If it produces no proofs, it gives no answer and is wrong. Nellie searches for up to $p$=10 proofs of max depth $d$=5 with a timeout of $t$=180 seconds per option.
Baselines
We evaluate Nellie against Entailer Tafjord et al. (2022), another system that produces entailment tree proofs via backward chaining. Entailer stops recurring not when it finds entailing facts from a corpus, but rather when the model believes that a subquery is true with high confidence. Its trees are thus not grounded in verified facts. We reimplemented their algorithm using their T5-11B-based model, reporting their configuration of a max tree depth (D) of 3 and minimum of 1 (i.e. $\geq 1$ decomposition). We also evaluate their ablated setting which recurs exactly once (D=1). This is a baseline that is neither grounded nor particularly interpretable, generating just a pair of statements. To isolate the impact of model size, we also evaluate an Entailer-3B model based on the same T5-3B model as Nellie’s DRG. [^8] We obtained the training data from Tafjord et al. (2022) and trained our model in consultation with the authors. We also compare against grounded xQA methods without logical structure (see Fig 2): PathNet Kundu et al. (2019), which constructs 2-hop chains by linking entities between facts, and three approaches that build graphs from facts using linear programming: TupleILP Khot et al. (2017), ExplanationLP Thayaparan et al. (2021), and Diff-Explainer Thayaparan et al. (2022). [^9] We show results from [Thayaparan et al. (2021) & 2022].
@@ 5.1 Results
Figure 6 shows the results of ablating Nellie’s two conditional generation modules and replacing them with vanilla generation. Removing both template- and retrieval-conditioned generation lowers Nellie’s performance in all circumstances, highlighting the empirical benefit of structured guidance. Ablating TCG ((-- TCG), and the drop from (-- RCG) to (-- Both)) reduces performance by 1-1.5 points. Ablating RCG (Nellie vs (-- RCG)) drops it by 3-4.
@@ 5.2 Domain Generalization and Knowledge Scaling
While Nellie was trained on datasets centered around WorldTree, we show that it can perform in another domain by altering the knowledge over which it reasons. We consider OpenBookQA Mihaylov et al. (2018), a dataset that requires reasoning over facts not included in WorldTree; each question is associated with one WorldTree science fact (“metals conduct electricity”), but also one fact from a separate pool of common knowledge (“a suit of armor is made of metal”).
Figure 6: Effect of ablating one or both of rule-conditioned (RCG) and template-conditioned generation (TCG).
@@
As with a classical expert system, we have designed Nellie to be “improvable merely by making statements to it” McCarthy (1959). This suggests that as we increase the provided knowledge store, we should expect Nellie to reason more effectively. Because WorldTree does not cover all the knowledge required to answer OBQA questions, we test whether Nellie performs better on them when we add the common knowledge to its retrieval index. As the common knowledge annotations are one fact per question and are not designed specifically for entailment, it is unlikely a priori that a fully grounded entailment tree can be created using the available knowledge alone, making this a challenging task.